7) Your yearly loan repayments on a $250,000 loan over ten years are $34,776.17. If the interest rate is 6.5% p.a., what is the amount of principal repaid in the first year?


a. $231,473.83

b. $16,250.00

c. $215,223.83

d. $18,526.17
Answer #1

Interest for first year=250,000*6.5%=$16250

Hence principal repayment for first year

=(34776.17-16250)

=$18526.17